# Scissorhands: Scrub Data Influence via Connection Sensitivity in Networks

### v2

@@ Abstract
Machine unlearning has become a pivotal task to erase the influence of data from a trained model. It adheres to recent data regulation standards and enhances the privacy and security of machine learning applications. In this work, we present a new machine unlearning approach Scissorhands. Initially, Scissorhands identifies the most pertinent parameters in the given model relative to the forgetting data via connection sensitivity. By reinitializing the most influential top-$k$ percent of these parameters, a trimmed model for erasing the influence of the forgetting data is obtained. Subsequently, Scissorhands fine-tunes the trimmed model with a gradient projection-based approach, seeking parameters that preserve information on the remaining data while discarding information related to the forgetting data. Our experimental results, conducted across image classification and image generation tasks, demonstrate that Scissorhands, showcases competitive performance when compared to existing methods.
Warning:
@@ 2.1 Overview
In Scissorhands, as shown in Figure 1, we initially identify critical parameters w.r.t. the forgetting data via connection sensitivity analysis [30, 31]. We then reinitialize the top-$k$ percent of these key parameters, resulting in a trimmed model with parameters ${\bm{\theta}}_{t}$. This step aims to diminish the model’s memory of the forgetting data ${\mathcal{D}}_{f}$. Nevertheless, this process runs the risk of erasing valuable information from the data set ${\mathcal{D}}_{r}$ that we aim to retain.
@@
Figure 1: Overview of Scissorhands. Given a model with parameter ${\bm{\theta}}$ trained using ${\mathcal{D}}$, in the trimming process, key parameters w.r.t. the forgetting data ${\mathcal{D}}_{f}$ would be pinpointed via Equation 2, and top-$k$of these parameters would be re-initialized for erasing the influence of ${\mathcal{D}}_{f}$, resulting in the trimmed model. In the repairing process, a gradient projection-based method is adopted for restoring the trimmed model utility over the remaining data ${\mathcal{D}}_{r}$ while excluding information associated with ${\mathcal{D}}_{f}$.
@@ 2.2 Trimming
Detailed proofs can be found in Section 7. Equation 2 defines the sensitivity of parameter $j$ as the average (i.e., expectation) of a product, the gradient of the loss and the current value of the parameter. As such, $\texttt{s}_{j}$ encodes the loss sensitivity according to the parameter’s magnitude. Intuitively, a parameter with a large value that significantly affects the loss, as captured by the gradient, is considered more influential or salient. The form in Equation 2 offers several benefits, most importantly, the ability to calculate the sensitivity for all parameters with just a single sweep through the dataset.
@@
Remark 1.
@@ 2.3 Repairing
Then, the optimal descent direction ${\bm{g}}$ is given from the solution $v^{*}$ in Section 2.3 as ${\bm{g}}={\bm{g}}_{o}-v^{*}{\bm{g}}_{f}$. Detailed proofs can be found in  Section 7. By doing so, we ensure a balance between unlearning and retaining utility.
Remark 3.
@@ 2.4 Algorithm Description
Algorithm 1 The procedure of Scissorhands.
1:procedure Trimming(${\bm{\theta}},{\mathcal{D}}_{f}$)
3:     // Compute connection sensitivity
4:     Get $\texttt{s}_{j}({\bm{\theta}};{\bm{X}},{\bm{y}}),\hskip 1.99997pt\forall j\in[1,d]$ using Eq. 2.
5:     // Re-initialization
6:     Re-initialize parameters according to the
7:       top-$k$value of $\texttt{s}_{j}$.
8:     return ${\bm{\theta}}_{t}$
9:end procedure
10:
11:procedure Repairing(${\bm{\theta}}_{t},{\mathcal{D}}_{f},{\mathcal{D}}_{r}$)
12:     ${\bm{\theta}}^{0}={\bm{\theta}}_{t}$.
13:     for $e=0$ to $E-1$ do
14:         Get $L({\bm{\theta}},{\mathcal{D}}_{r},{\mathcal{D}}_{f})$ (cf. Eq. 3).
15:         Get ${\bm{g}}_{o}$ and ${\bm{g}}_{f}$.
16:         // Get the optimal direction
17:         if $\langle{\bm{g}}_{o},{\bm{g}}_{f}\rangle>0$ then
18:              Compute $v^{*}$ (cf. Sec. 2.3).
19:              ${\bm{g}}={\bm{g}}_{o}-v^{*}{\bm{g}}_{f}$.
20:              ${\bm{g}}_{o}\leftarrow{\bm{g}}$.
21:         end if
22:         ${\bm{\theta}}^{e+1}\leftarrow{\bm{\theta}}^{e}-\eta{\bm{g}}_{o}$.
23:     end for
24:     return ${\bm{\theta}}^{*}={\bm{\theta}}^{E}$.
25:end procedure
Algorithm 1 describes the procedure of Scissorhands in detail. We first identify the important parameters w.r.t. the forgetting data in single-shot via connection sensitivity in networks, then re-initialize these parameters for erasing the influence of the forgetting data ${\mathcal{D}}_{f}$ in the given model. To repair the performance of the trimmed model on the remaining data ${\mathcal{D}}_{r}$, while ensuring the exclusion of information associated with the forgetting data ${\mathcal{D}}_{f}$, we employ the gradient projection-based optimization algorithm. By doing so, we can scrub the influence of the forgetting data while retaining model utility on the remaining data.
@@ 3 Related Work
The development of efficient machine unlearning methods [43, 26, 3, 16, 2, 40, 55, 21, 17, 35, 46, 25, 48, 38] has gained prominence. Applications of machine unlearning span various domains, including regression tasks [48], federated learning [33, 32, 22, 54, 52], graph neural networks [4, 7], and diffusion models [14, 57, 24, 15, 11], as well as scenarios where training data are not available [49, 8].
@@ 4 Experimental Evaluation
In this section, we illustrate how Scissorhands effectively eliminates the data influence in the models. For sample-wise forgetting, where the forgetting data shares the same distribution as the training data, we evaluate Scissorhands on SVHN [37], CIFAR-10 and CIFAR-100 [28]. We further extend our evaluation on CelebAMask-HQ [29] where targeting to remove the entirety of specific identities. Additionally, our experimentation encompasses the open-source text-to-image model, Stable Diffusion v1.4 [42], which is conditioned on CLIP text embeddings through the cross-attention mechanism. Further detailed experimental setups and additional results can be found in Sec. 8.
Baselines.
@@
We primarily compare against the following standard baselines that are frequently employed in machine unlearning, alongside recently proposed SOTA methods: (i) Retrain: models obtained by only retraining from scratch on ${\mathcal{D}}_{r}$. This will provide the performance of the oracle for us. (ii) Fine-tuning (FT) [53]: models that are fine-tuned on ${\mathcal{D}}_{r}$, i.e., taking advantage from catastrophic forgetting in neural networks to unlearn. (iii) Gradient ascent (GA) [50]: gradient ascent on ${\mathcal{D}}_{f}$. This will provide a simple way of unlearning by making the performance of the model worse on ${\mathcal{D}}_{f}$. (iv) Influence unlearning (IU) [27]: utilizes the influence function method to estimate the change in model parameters when transitioning from the unscrubbed model to the retrained model. (v) Boundary shrink (BS) [5] and (vi) Boundary expanding (BE) [5]: shift the decision boundary of the original model to imitate that of the retrained model. (vii) $\ell_{1}$-sparse [25]: fine-tuning models on ${\mathcal{D}}_{r}$ with $\ell_{1}$-norm sparse regularization. (viii) Saliency unlearning (SalUn) [11]: adopt weight saliency and random labeling for unlearning.
Metrics.
@@
To assess the unlearning algorithms, we employ several metrics: (i) Accuracy: the accuracy of the model on ${\mathcal{D}}_{f}$ (denoted as $\text{Acc}_{{\mathcal{D}}_{f}}$), ${\mathcal{D}}_{r}$ (denoted as $\text{Acc}_{{\mathcal{D}}_{r}}$) and the test set (denoted as $\text{Acc}_{{\mathcal{D}}_{t}}$). It provides insight into how well the model performs after undergoing the unlearning process. (ii) Membership inference attack (MIA): a standard metric for verifying the unlearning effect. A classifier $\phi$ is trained using both training data (marked with a label of 1, indicating data points included in training) and test data (marked with a label of 0, indicating data points not seen during training), and subsequently assessed on ${\mathcal{D}}_{f}$. An effective unlearning method should make it challenging to infer whether a particular sample was part of the training set, mitigating potential privacy breaches. In this context, we define MIA as $P_{\phi}({\bm{y}}=0|{\bm{x}}_{f})$, representing the probability assigned by the classifier $\phi$ that a given sample ${\bm{x}}_{f}$ was not included in the training data. (iii)Avg. Gap [11]: the average of the performance gaps measured in accuracy-related metrics: Avg. Gap $=(|\text{Acc}_{{\mathcal{D}}_{t}}-\text{Acc}^{\ast}_{{\mathcal{D}}_{t}}|+|\text{Acc}_{{\mathcal{D}}_{f}}-\text{Acc}^{\ast}_{{\mathcal{D}}_{f}}|+|\text{Acc}_{{\mathcal{D}}_{r}}-\text{Acc}^{\ast}_{{\mathcal{D}}_{r}}|+|\text{MIA}-\text{MIA}^{\ast}|)/4$. where $\text{Acc}^{\ast}_{{\mathcal{D}}_{t}},\text{Acc}^{\ast}_{{\mathcal{D}}_{f}},\text{Acc}^{\ast}_{{\mathcal{D}}_{r}}$ and $\text{MIA}^{\ast}$ are metric values of the retrained model. A better performance of an unlearning method corresponds to a lower performance gap with retraining, measuring how close the unlearned model is to the retrained model. (iv) Run-time efficiency (RTE) [25]: evaluate the computational efficiency of the unlearning method. Specifically, we have $\text{RTE}={\text{T}}/{\text{T}_{r}}$, where T and $\text{T}_{r}$ present the time for the unlearning methods to get the scrubbed models and the time for retraining from scratch with ${\mathcal{D}}_{r}$, respectively. An efficient unlearning method should aim for minimal computational overhead compared to the baseline. (v) Relearn time [17, 8, 48]: the number of epochs to fine-tune the scrubbed models for regaining performance on ${\mathcal{D}}_{f}$ (i.e., reach the original models’ accuracy on ${\mathcal{D}}_{f}$). The larger the number, the more effective the erasure.
@@ 4.1 Results on Classification Task
Table 1: Quantitative results for forgetting 10% data on the SVHN, CIFAR-10 and CIFAR-100 datasets.

### v3

@@ Abstract
Machine unlearning has become a pivotal task to erase the influence of data from a trained model. It adheres to recent data regulation standards and enhances the privacy and security of machine learning applications. In this work, we present a new machine unlearning approach Scissorhands. Initially, Scissorhands identifies the most pertinent parameters in the given model relative to the forgetting data via connection sensitivity. By reinitializing the most influential top-$k$ percent of these parameters, a trimmed model for erasing the influence of the forgetting data is obtained. Subsequently, Scissorhands fine-tunes the trimmed model with a gradient projection-based approach, seeking parameters that preserve information on the remaining data while discarding information related to the forgetting data. Our experimental results, conducted across image classification and image generation tasks, demonstrate that Scissorhands, showcases competitive performance when compared to existing methods. Source code is available at https://github.com/JingWu321/Scissorhands.
0.0.1 Warning:
@@
Keywords:
Machine unlearning Connection sensitivity Diffusion model
@@ 2 Methodology
In Scissorhands, we initially identify critical parameters w.r.t. the forgetting data via connection sensitivity analysis [33, 32]. We then reinitialize the top-$k$ percent of these key parameters, resulting in a trimmed model with parameters ${\bm{\theta}}_{t}$. This step aims to diminish the model’s memory of the forgetting data ${\mathcal{D}}_{f}$. Nevertheless, this process runs the risk of erasing valuable information from the data set ${\mathcal{D}}_{r}$ that we aim to retain.
@@ 2.1 Trimming
Detailed proofs can be found in Appendix A. Equation 2 defines the sensitivity of parameter $j$ as the average (i.e., expectation) of a product, the gradient of the loss and the current value of the parameter. As such, $\texttt{s}_{j}$ encodes the loss sensitivity according to the parameter’s magnitude. Intuitively, a parameter with a large value that significantly affects the loss, as captured by the gradient, is considered more influential or salient. The form in Equation 2 offers several benefits, most importantly, the ability to calculate the sensitivity for all parameters with just a single sweep through the dataset.
@@
Remark 1
In principle, Scissorhands can benefit from any algorithms that can identify important connections w.r.t. $\mathcal{D}_{f}$. The main reason behind choosing SNIP is its single-shot nature. As presented in the study [12], the gradient of loss w.r.t. the model parameters over the forgetting data can help identify important parameters w.r.t. $\mathcal{D}_{f}$ and hence help with erasing.
Remark 2
@@ 2.2 Repairing
Algorithm 1 The procedure of Scissorhands.
1:procedure Trimming(${\bm{\theta}},{\mathcal{D}}_{f}$)
3:     // Compute connection sensitivity
4:     Get $\texttt{s}_{j}({\bm{\theta}};{\bm{X}},{\bm{y}}),\hskip 1.99997pt\forall j\in[1,d]$
5:       using Eq. 2.
6:     // Re-initialization
7:     Re-initialize parameters according
8:       to the top-$k$value of $\texttt{s}_{j}$.
9:     return ${\bm{\theta}}_{t}$
10:end procedure
11:
12:procedure Repairing(${\bm{\theta}}_{t},{\mathcal{D}}_{f},{\mathcal{D}}_{r}$)
13:     ${\bm{\theta}}^{0}={\bm{\theta}}_{t}$.
14:     for $e=0$ to $E-1$ do
15:         Get $L({\bm{\theta}},{\mathcal{D}}_{r},{\mathcal{D}}_{f})$ (cf. Eq. 3).
16:         Get ${\bm{g}}_{o}$ and ${\bm{g}}_{f}$.
17:         // Get the optimal direction
18:         if $\langle{\bm{g}}_{o},{\bm{g}}_{f}\rangle>0$ then
19:              Compute $v^{*}$ (cf. Sec. 2.2).
20:              ${\bm{g}}={\bm{g}}_{o}-v^{*}{\bm{g}}_{f}$.
21:              ${\bm{g}}_{o}\leftarrow{\bm{g}}$.
22:         end if
23:         ${\bm{\theta}}^{e+1}\leftarrow{\bm{\theta}}^{e}-\eta{\bm{g}}_{o}$.
24:     end for
25:     return ${\bm{\theta}}^{*}={\bm{\theta}}^{E}$.
26:end procedure
@@
Then, the optimal descent direction ${\bm{g}}$ is given from the solution $v^{*}$ in Section 2.2 as ${\bm{g}}={\bm{g}}_{o}-v^{*}{\bm{g}}_{f}$. Detailed proofs can be found in Appendix A. By doing so, we ensure a balance between unlearning and retaining utility.
Remark 4
@@ 2.3 Algorithm Description
Algorithm 1 describes the procedure of Scissorhands in detail. We first identify the important parameters w.r.t. the forgetting data in single-shot via connection sensitivity in networks, then re-initialize these parameters for erasing the influence of the forgetting data ${\mathcal{D}}_{f}$ in the given model. To repair the performance of the trimmed model on the remaining data ${\mathcal{D}}_{r}$, while ensuring the exclusion of information associated with the forgetting data ${\mathcal{D}}_{f}$, we employ the gradient projection-based optimization algorithm. By doing so, we can scrub the influence of the forgetting data while retaining model utility on the remaining data. The loss depicted in Equation 3 can be viewed as a scalarization of a Multi-Objective Optimization (MOO) problem, namely minimizing $\big{(}\mathcal{L}(\theta;\mathcal{D}_{r}),-\mathcal{L}(\theta;\mathcal{D}_{f})\big{)}^{\top}$. A common issue in MOO is the gradient conflict, as optimizing one objective could hinder another one. Our proposal on gradient projection is designed to address and mitigate the gradient conflict. Empirically (see Table 3), we studied the effect of removing gradient projection.
@@ 3 Related Work
The development of efficient machine unlearning methods [45, 28, 3, 17, 2, 42, 58, 23, 20, 37, 48, 27, 51, 40, 18, 11, 61] has gained prominence. Applications of machine unlearning span various domains, including regression tasks [51], federated learning [35, 34, 24, 56, 54], graph neural networks [5, 7], and diffusion models [15, 60, 26, 16, 12, 57, 62], as well as scenarios where training data are not available [50, 8].
@@ 4 Experimental Evaluation
In this section, we illustrate how Scissorhands effectively eliminates the data influence in the models. For sample-wise forgetting, where the forgetting data shares the same distribution as the training data, we evaluate Scissorhands on SVHN [39], CIFAR-10 and CIFAR-100 [30]. We further extend our evaluation on CelebAMask-HQ [31] where targeting to remove the entirety of specific identities. Additionally, our experimentation encompasses the open-source text-to-image model, Stable Diffusion v1.4 [44], which is conditioned on CLIP text embeddings through the cross-attention mechanism. Further detailed experimental setups and additional results can be found in Appendix B.
@@
Baselines. We primarily compare against the following standard baselines that are frequently employed in machine unlearning, alongside recently proposed SOTA methods: (i) Retrain: models obtained by only retraining from scratch on ${\mathcal{D}}_{r}$. This will provide the performance of the oracle for us. (ii) Fine-tuning (FT) [55]: models that are fine-tuned on ${\mathcal{D}}_{r}$, i.e., taking advantage from catastrophic forgetting in neural networks to unlearn. (iii) Gradient ascent (GA) [52]: gradient ascent on ${\mathcal{D}}_{f}$. This will provide a simple way of unlearning by making the performance of the model worse on ${\mathcal{D}}_{f}$. (iv) Influence unlearning (IU) [29]: utilizes the influence function method to estimate the change in model parameters when transitioning from the unscrubbed model to the retrained model. (v) Boundary shrink (BS) [4] and (vi) Boundary expanding (BE) [4]: shift the decision boundary of the original model to imitate that of the retrained model. (vii) $\ell_{1}$-sparse [27]: fine-tuning models on ${\mathcal{D}}_{r}$ with $\ell_{1}$-norm sparse regularization. (viii) Saliency unlearning (SalUn) [12]: adopt weight saliency and random labeling for unlearning.
@@
Metrics. To assess the unlearning algorithms, we employ several metrics: (i) Accuracy: the accuracy of the model on ${\mathcal{D}}_{f}$ (denoted as $\text{Acc}_{{\mathcal{D}}_{f}}$), ${\mathcal{D}}_{r}$ (denoted as $\text{Acc}_{{\mathcal{D}}_{r}}$) and the test set (denoted as $\text{Acc}_{{\mathcal{D}}_{t}}$). It provides insight into how well the model performs after undergoing the unlearning process. (ii) Membership inference attack (MIA): a standard metric for verifying the unlearning effect. A classifier $\phi$ is trained using both training data (marked with a label of 1, indicating data points included in training) and test data (marked with a label of 0, indicating data points not seen during training), and subsequently assessed on ${\mathcal{D}}_{f}$. An effective unlearning method should make it challenging to infer whether a particular sample was part of the training set, mitigating potential privacy breaches. In this context, we define MIA as $P_{\phi}({\bm{y}}=0|{\bm{x}}_{f})$, representing the probability assigned by the classifier $\phi$ that a given sample ${\bm{x}}_{f}$ was not included in the training data. (iii)Avg. Gap [12]: the average of the performance gaps measured in accuracy-related metrics: Avg. Gap $=(|\text{Acc}_{{\mathcal{D}}_{t}}-\text{Acc}^{\ast}_{{\mathcal{D}}_{t}}|+|\text{Acc}_{{\mathcal{D}}_{f}}-\text{Acc}^{\ast}_{{\mathcal{D}}_{f}}|+|\text{Acc}_{{\mathcal{D}}_{r}}-\text{Acc}^{\ast}_{{\mathcal{D}}_{r}}|+|\text{MIA}-\text{MIA}^{\ast}|)/4$. where $\text{Acc}^{\ast}_{{\mathcal{D}}_{t}},\text{Acc}^{\ast}_{{\mathcal{D}}_{f}},\text{Acc}^{\ast}_{{\mathcal{D}}_{r}}$ and $\text{MIA}^{\ast}$ are metric values of the retrained model. A better performance of an unlearning method corresponds to a lower performance gap with retraining, measuring how close the unlearned model is to the retrained model. (iv) Run-time efficiency (RTE) [27]: evaluate the computational efficiency of the unlearning method. Specifically, we have $\text{RTE}={\text{T}}/{\text{T}_{r}}$, where T and $\text{T}_{r}$ present the time for the unlearning methods to get the scrubbed models and the time for retraining from scratch with ${\mathcal{D}}_{r}$, respectively. An efficient unlearning method should aim for minimal computational overhead compared to the baseline. (v) Relearn time [20, 8, 51]: the number of epochs to fine-tune the scrubbed models for regaining performance on ${\mathcal{D}}_{f}$ (i.e., reach the original models’ accuracy on ${\mathcal{D}}_{f}$).
@@ 4.1 Results on Classification Task
Table 1: Quantitative results for forgetting 10% data.